OpenStudy (anonymous):

check all that apply. If csc theta = 13/12 then: a. tan theta = 12/5 b. sin theta = 12/13 c. cos theta = 12/13 d. sec theta = 12/13

OpenStudy (jdoe0001):

\(\bf csc(\theta)=\cfrac{13}{12}\implies \cfrac{hypotenuse}{opposite}\implies \cfrac{c=13}{b=12}\\ \quad \\ \textit{using the pythagorean theorem }c^2={\color{blue}{ a}}^2+b^2\implies \sqrt{c^2-b^2}={\color{blue}{ a}}\) once you find "a", or adjacent side, check against your choices

OpenStudy (jdoe0001):

keeping in mind -> http://www.mathplanet.com/images/math/codecogs_cbc49307.gif

OpenStudy (anonymous):

a = 25

OpenStudy (anonymous):

how would i check?

OpenStudy (jdoe0001):

a = 25?

OpenStudy (anonymous):

I think so but how would i find out what else works?

OpenStudy (jdoe0001):

\(\bf c^2={\color{blue}{ a}}^2+b^2\implies \sqrt{c^2-b^2}={\color{blue}{ a}}\implies \sqrt{13^2-12^2}={\color{blue}{ a}}\) you plug in your a, b and c, in the trig identities given, tangent, sine, cosine, secant see if they match with the values for "a", "b", and "c" that you have

OpenStudy (anonymous):

So would the answer be B, and C?

OpenStudy (jdoe0001):

\(\bf \sqrt{13^2-12^2}={\color{blue}{ a}}\implies \sqrt{169-144}=a\implies \sqrt{25}=a\implies 5=a\\ \quad \\ \textit{let's test the first choice, say }tan(\theta)\\ \quad \\ tan(\theta)=\cfrac{b}{a}\implies \cfrac{13}{5}\) there's no 1 answer all may be correct or some of them only so you'd have to check all

OpenStudy (jdoe0001):

hmmm dohhh I meant 12.. lemme fix that \(\bf \sqrt{13^2-12^2}={\color{blue}{ a}}\implies \sqrt{169-144}=a\implies \sqrt{25}=a\implies 5=a\\ \quad \\ \textit{let's test the first choice, say }tan(\theta)\\ \quad \\ tan(\theta)=\cfrac{b}{a}\implies \cfrac{12}{5}\)

OpenStudy (mertsj):



OpenStudy (jdoe0001):

so as you can see, the 1st checks out

OpenStudy (anonymous):

Oh ok, so it would end up being A and B right

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

alright, thanks

OpenStudy (jdoe0001):

yw